# Prompts to Summaries: Zero-Shot Language-Guided Video Summarization

### v1

@@ 3.4 User guided scoring
this preference. Scenes that closely match or contradict the user’s intent should be scored accordingly,
reflecting their relevance or
@@
If the scene is not clearly related to this preference, assign a score based on the default scale and
criteria below.
3.5 Frame level scoring
To obtain frame-level importance scores, we follow a three-stage procedure. First, each frame initially inherits the importance score of its parent scene, as shown in fig. 4c. Then, a temporal smoothing function is applied around scene boundaries to avoid abrupt changes and create a gradual transition between adjacent scenes. After that, within every scene, we compute a representativeness weight for each frame (how well it reflects that scene). Finally, this weight is then fused with the smoothed scene score, yielding the final per-frame importance value for the entire video.
Goal and approach: After establishing scene boundaries (Figure 4b) and computing the score of each scene (Figure 4c), a normalization step is applied to the scores of the scenes to emphasize relative differences in importance as perceived by the LLMs. Since the scores are not derived from a binary scale (important or not important), but rather fall along a spectrum of relevance (1-100), normalization helps highlight subtle distinctions and amplifying the contrast between different scenes, ensuring that highly important scenes are more prominently reflected while less relevant scenes are de-prioritized. See Section 4.1.4 for the exact used normalization procedure. Now, each frame within the scene is initially assigned the same normalized score as the scene to which it belongs (Figure 4d).
3.5.2 Frame weighting with in a scene
Goal: To go beyond our scene-level importance scores—which already summarize contextual, semantic, and temporal cues—we now assign frame-level weights. Each weight quantifies how well a given frame describes its entire scene, letting us pinpoint the most representative and informative moments within that scene; i.e., assessing every frame a representativeness score within the scene.
Approach: First, we note that Algorithm 4 in the appendix provides the full details for this step. We first segment each scene into non-overlapping segments each of size $W$. For each segment, we define and compute two key properties to assess its contribution to the overall summary: consistency and uniqueness. These two scores aim to balance stability and novelty in the selected content. To compute these properties, we begin by employing KMeans [37, 38] for clustering the frame embeddings of each scene, to reveal the underlying structure and visual patterns. The optimal number of cluster for each scene, donated as $K^{*}$, is selected dynamically by evaluating a predefined range of candidate values; For each candidate value $K_{i}\in\{K_{min},K_{min}+\Delta{k},\dots,K_{max}\}$, we compute Within-Cluster Sum of Squares (WCSS)  [39], donated, we then apply the Elbow method [40] to the WCSS curve to select $K^{*}$, the value at which the rate of decrease in WCSS significantly slows down—identified as the point with the greatest change in the slope of the WCSS curve.
Consistency Reflects the internal coherence of the segment. It quantifies how uniformly the frames within a segment belong to the same visual pattern, as indicated by their cluster labels. A segment with high consistency suggests a stable and semantically meaningful portion of the video. Formally, given a segment $S$ of $W$ consecutive frames with cluster labels $\{l_{t+1},l_{t+2},\dots,l_{t+W}\}$, the consistency score is defined as :
where $c$ ranges over all cluster labels present in the segment $S$ of size $W$.
Uniqueness (or dissimilarity) measures how distinct the segment is with respect to internal variation in visual features. It captures how much the frame embeddings in the segment deviate from the segment’s mean, which helps highlight rare or visually rich content. Given the frame embeddings $\{f_{t+1},f_{t+2},\dots,f_{t+W}\}$ in segment $S$, we compute the mean embedding $\bar{\mathbf{f}}$ and define uniqueness as:
$\text{Uniqueness}(S)=\frac{1}{W}\sum_{i=1}^{W}\left\|\mathbf{f}_{t+i}-\bar{\mathbf{f}}\right\|_{2},\bar{\mathbf{f}}=\frac{1}{W}\sum_{i=1}^{W}\mathbf{f}_{t+i},\quad$ (6)
Finally, we combine these two factors — consistency and uniqueness — into a unified score to assign a weight to each segment.
$Weight(S)=\sigma\cdot Consistency(S)+(1-\sigma)\cdot Uniqueness(S)$ (7)
The weighting of these factors is governed by the parameter $\sigma$, which allows us to adjust the relative importance of consistency versus uniqueness. All frames within a given segment inherit the segments computed weight. Furthermore, the frame consistency weight $\sigma$ and segment duration $W$ are dynamically determined according to the video duration, donated as $T$,as detailed in Equation 8.
$F_{S}(\text{Video})=\begin{cases}\{\sigma{=}0.1,\;W{=}1\,\text{s}\}&\text{if }T>5S\\ \{\sigma{=}1.0,\;W{=}1\,\text{s}\}&\text{if }T>S\\ \{\sigma{=}0.3,\;W{=}3\,\text{s}\}&\text{otherwise}\end{cases}$ (8)
Where $S$ represents a short video duration threshold, $F_{S}$ selects the optimal frame consistency weight $\sigma$ and segment duration $W$ based on the video’s length. We categorize videos into three classes: (i) Long videos — duration > $5S$ (ii) Intermediate videos — $S<$ duration $\leq 5S$ (iii) Short videos — duration $\leq S$
We set $S=1.8$ minutes (approximately 100 seconds), meaning a video is considered long if its duration exceeds 9 minutes. Additional details behind the design choices are provided in Section  4.4.
Input: Video frames $F_{Video}=\{f_{1},f_{2},\dots,f_{T}\}$
Optional : $User\;Query=Q_{user}$
Output: Frame-level importance scores
Initialization: $\tilde{\mathcal{S}}\leftarrow\emptyset$ ;
//Detect scenes boundaries(Section  3.1)
@@ 3.5.2 Frame weighting with in a scene
1. $\mathcal{B}\leftarrow Scene\_Boundry\_Detection(F_{Video})$ //Generate scenes and full video descriptions(Section 3.2
2. $Scenes_{Desc},Video_{Desc}\leftarrow Descriptions(\mathcal{B},F_{Video})$ //Score scenes using LLM( 3.3)
3. $\mathcal{S}\leftarrow Scene\_Scores(Scenes_{Desc},Video_{Desc},Q_{user})$ //Normalize scenes scores and smooth frame-level scores(Section  3.5.1
4. ${F}_{smooth}\leftarrow Scene\_Scores\_Smoothing(\mathcal{S},\mathcal{B},F_{norm})$ //Dynamically select segment duration and frame consistency weight(Equation 8)
5. $W,\sigma\leftarrow F_{s}(F_{Video})$ //Compute frame weights(Section 3.5.2)
6. ${F}_{weights}\leftarrow Frame\_Weighting(F,\mathcal{B},W,\sigma)$
7. foreach  scene $i=1,\dots,N$  do
       Get scene boundaries $\mathcal{B}_{i}=[start_{i},end_{i}]$ ;
       $\tilde{\mathcal{S}}_{i}\leftarrow\emptyset$ ;
      foreach  frame $t$ in $\mathcal{B}_{i}$  do
             //Compute final frame score(Equation 7)
             $\tilde{f_{t}}\leftarrow{F}_{smooth}[t]\times{F}_{weights}[t]$ ;
             Append $\tilde{f_{t}}$ to $\tilde{\mathcal{S}}_{i}$ ;
      Normalize within scene
       $\tilde{\mathcal{S}}_{i}\leftarrow F_{norm}(\tilde{\mathcal{S}}_{i})$;
       Append $\tilde{\mathcal{S}}_{i}$ to $\tilde{\mathcal{S}}$ ;
return Frame-level importance score $\tilde{\mathcal{S}}$
@@ 6. ${F}_{weights}\leftarrow Frame\_Weighting(F,\mathcal{B},W,\sigma)$
3.5.3 Frame-level importance scoring
Goal: Fusing scene importance and frame weights to create per-frame importance.
Approach: The final step calculates frame-level importance scores, which are the output of our algorithm; providing the importance each frame within the video. Given the smoothed per-frame score (from Section 3.5.1) and per-frame wights (as in Section 3.5.2), for each frame within a scene, we compute the final frame scores as the product of the smoothed scores, $F_{smooth}$ and the frame weights, $F_{weights}$. These frame-level scores are then normalized, by applying the same normalization function $F_{norm}$, within each scene. We apply normalization of frame scores within each scene-rather than the entire video, to ensure that the frame’s importance is evaluated based on the context of their own scene and not based on the context of the overall video. This ensues that every scene within the video contributes fairly to the final summary, based on its own internal dynamics and relative frame importance.
The result is a set of frame-level importance scores (Figure 4e) $\tilde{\mathcal{S}}$, which represent the contribution of each frame to the overall video summary. This process, as shown in Algorithm 1, ensures that the frame-level importance scores take into account both the scene’s intrinsic importance, the temporal and contextual relationships between frames.
4 Experiments and results
4.1 Standard video summarization
4.1.1 Datasets
We utilize two widely used datasets: SumMe [25] and TVSum [26]. The SumMe dataset consist of 25 videos, ranging from 1 to 6 minutes in duration, comprising of a various range of content including holidays, events, and sports; captured from both first-person and third-person perspectives. Each video has been annotated by 15-18 users, who provided key-fragments. Consequently, each video is accompanied by multiple user-generated summaries of varying lengths, typically ranging from 5% to 15% of the original video’s duration. The TVSum dataset consist of 50 videos, ranging from 1 to 61 minutes in duration. These videos belong to 10 different categories comprising of various activities like changing a vehicle tire, making a sandwich, and visiting an award wining sandwich shop, with 5 videos in each category, and were collected from YouTube. For annotation, each video was reviewed and rated by 20 users, who assigned frame-level importance scores on a scale from 1 (not important) to 5 (very important). [15]
4.1.2 Evaluation protocol
We utilize the classic keyshot evaluation method, introduced by [6] and widely adopted by state-of-the-art video summarization methods. This metric assesses the similarity between the generated summary (A), limited to less than 15% of the the original video duration, and the user provided summary (B). It compute the temporal overlap between the summaries using (P)recision and (R)ecall:
@@ 4.1.2 Evaluation protocol
where $|A|$ denotes the duration of the set $A$, measured in number of frames. Then we compute the F1-score (F), which represents the harmonic mean of (P)recision and (R)ecall, as:
F $=2\cdot\frac{P\cdot R}{P+R}$ (10)
In SumMe, annotations are provided as key-fragments, this formula is directly applied. However, in the TVSum, annotations are provided at the frame level, with translation from the frame level to the key fragment using methodologies presented in [6]. After obtaining key-fragment annotations for a given test video, we calculate the F1-score for the generated summary against each user summary for that video. The F1 scores for the TVSum dataset are averages as proposed inr [26], while for the SumMe dataset, the F1 score for a given test video is the maximum F1 score calculated for that video, as proposed in its paper [41]. To obtain the final F1 score for each data set, we calculate the mean F1 score in the test videos, and take the mean results in 5 different splits of generated video tests, following the methodology in  [6]. This score serves as a performance metric.
4.1.3 Design choices
For each dataset, we choose one of three normalization functions—Min-Max normalization, Exponential normalization, or their combination—for the normalization step described in Section 3.5.1, based on the dataset’s characteristics.
The used normalization for SumMe. Min-Max normalization linearly scales values between 0 and 1, preserving relative distances but remaining sensitive to outliers. This characteristic makes it particularly well-suited to the SumMe dataset, where misalignment between user-selected keyframes often occurs due to subjective differences in user preferences. As noted in [6], the evaluation protocol for SumMe selects the closest user summary by taking the maximum F1-score across all references. Consequently, we adopt Min-Max normalization to better align the computed importance scores with the variability inherent in human annotations.
The used normalization for TVSum. In contrast, Exponential normalization amplifies differences among higher values, making it particularly effective for highlighting prominent segments. This characteristic makes it especially well-suited to the TVSum dataset, where there is no misalignment between user-selected keyframes, and the user scores tend to be more uniformly assigned, as the evaluation protocol for TVSum is computed based on the average F1-score across all users. Therefore, we adopt Exponential normalization to better align the computed importance scores with the relative consistency (invariability) present in human annotations.
The combined approach aims to leverage the benefits of both techniques, balancing local contrast with global consistency in frame-level importance scores. To validate the correctness of our choices, we conducted additional experiments comparing different normalization methods across multiple variations on both datasets. Further details are provided in the Section  4.4.
Table 1: Performance evaluation of out model on TVSum and SumMe test sets (Precision, Recall, F1 Score)
Method Dataset SumMe TVSum (P)recision 55.77 62.223 (R)ecall 57.96 62.224 F1-Score 56.73 62.216
4.1.4 Results
Precision, recall, and F1 score. We evaluated our model’s precision, recall, and F1 score on the test sets of SumMe and TVSum to provide a fare comparison to other methods (no training is needed in our case). First, Table 1 reports the results of our method; the close alignment between precision, recall, and F1 across both datasets suggests that the model reliably selects relevant segments while maintaining strong coverage of human annotations. Notably, SumMe exhibits slightly higher precision and recall margins than TVSum, likely due to greater variability in its human-labeled summaries as mentioned before. With our conclusive scores of 56.73% on the SumMe dataset and 62.21% on the TVSum dataset in Table 1, these outcomes will be used for comparative analysis with other methods.
Table 2: Performance on TVSum and SumMe.Every baseline—supervised or “unsupervised”—relies on a training stage. Our pipeline is truly zero-shot: it needs no tuning, no access to ground-truth labels, and no prior knowledge of the datasets. Despite that handicap, it sets a new state-of-the-art among unsupervised methods and matches the best supervised models.
Method name Doesn’t require training data Text-queryable Dataset

### v2

@@ A.1 LLM prompt
  this preference. Scenes that closely match or contradict the user’s
  intent should be scored accordingly, reflecting their relevance or
@@
  If the scene is not clearly related to this preference, assign a
  score based on the default scale and criteria below.
Assign an importance score on a scale of 1 to 100, based on how crucial it is to the overall video. The scale is defined as follows:
* 1-20: Minimally important
 (contributes very little to the overall theme or message)
* 21-40: Somewhat important
 (offers limited context or details that support the main theme)
* 41-60: Moderately important
 (provides useful context or details that support the main theme)
* 61-80: Quite important
 (adds significant context or detail that enhances understanding of the main theme)
* 81-100: Highly important
 (crucial to understanding or conveying the main message of the video)
Video Description: <<video description>>
Scene Description: <<scene description>>"
A.2 Our algorithms
Algorithm 1 outlines the process for computing scene-level importance scores using a large language model (LLM). Each scene description is paired with the full video description and inserted into a prompt template. The resulting query (Section A.1) is then evaluated by the LLM to estimate how critical that scene is to the overall video, as described in details in Section  3.3
Algorithm 2 describes the scene score smoothing procedure. The method first normalizes scene-level scores and assigns each frame within a scene the scene’s score. To ensure smooth transitions between consecutive scenes, the algorithm linearly interpolates frame scores over the transition region defined by the midpoints between scenes, using a cosine-based weighting function. This smoothing step helps reduce abrupt changes in frame-level importance and enhances temporal coherence, as discussed in Section 3.4.1.
Algorithm 3 details the frame weighting procedure within each scene. The algorithm extracts frame embeddings and determines the optimal number of clusters using the Elbow method based on within-cluster sum of squares (WCSS). Frames are grouped into segments of fixed size $W$, and each segment is scored by computing its consistency and uniqueness measures, weighted by parameter $\sigma$. These segment-level scores are then assigned uniformly to the frames within each segment, providing refined frame-level weights that capture both local coherence and distinctiveness, as detailed in Section 3.4.2.
Algorithm 1 Scene-Level Importance Scores (Section 3.3)
@@ A.2 Our algorithms
1:Input: Scene descriptions $\mathcal{D}_{scenes}$, full video description $\mathcal{D}_{video}$, LLM prompt template $\mathcal{P}$
2:Output: Scene-level importance scores $\mathcal{S}$
3:Initialize scene importance scores list $\mathcal{S}\leftarrow\emptyset$
4:for $d_{i}\in\mathcal{D}_{scenes}$ do
5:  Construct query $q_{i}\leftarrow\text{Format}(\mathcal{P},d_{i},\mathcal{D}_{video})$
6:  Send $q_{i}$ to LLM and receive scene score $s_{i}$
7:  Append $s_{i}$ to $\mathcal{S}$
8:end for
9:return $\mathcal{S}$
Algorithm 2 Scene Scores Smoothing (Section 3.4.1)
1:Input: Scene scores $\{s_{1},\dots,s_{N}\}$, scene boundaries $\{\mathcal{B}_{1},\dots,\mathcal{B}_{N}\}$, normalization $F_{norm}$
2:Output: Frame-level scores $\{f_{1},\dots,f_{T}\}$
3:Initialize scene scores $\mathcal{S}\leftarrow F_{norm}(S)$
4:for $i=1,\dots,N$ do
5:  Get scene score $s_{i}$
6:  Get scene boundaries $\mathcal{B}_{i}=[start_{i},end_{i}]$
7:  for $t\in\mathcal{B}_{i}$ do
8:   Assign $f_{t}=s_{i}$
9:  end for
10:end for
11:for $i=1,\dots,N-1$ do
12:  $mid_{i}=\frac{start_{i}+end_{i}}{2},\quad mid_{i+1}=\frac{start_{i+1}+end_{i+1}}{2}$
13:  $L=mid_{i+1}-mid_{i}$
14:  for $t=mid_{i},\dots,mid_{i+1}$ do
15:   $p=\frac{t-mid_{i}}{L}$
16:   $w=\frac{1-\cos(\pi\cdot p)}{2}$
17:   $f_{t}=(1-w)\cdot s_{i}+w\cdot s_{i+1}$
18:  end for
19:end for
20:return $\{f_{1},\dots,f_{T}\}$
Algorithm 3 Frame Weighting (Section 3.4.2)
1:Input: Frames $\{f_{1},\dots,f_{T}\}$, scene boundaries $\{\mathcal{B}_{1},\dots,\mathcal{B}_{N}\}$, segment size $W$, parameter $\sigma$
2:Output: Frame weights $\{w_{1},\dots,w_{T}\}$
3:for $i=1,\dots,N$ do
4:  $E_{\mathcal{B}_{i}}=\{e_{t}\mid t\in\mathcal{B}_{i}\}$
5:  $K^{*}=\arg\min_{K_{i}\in[K_{\min},K_{\max}]}WCSS(E_{\mathcal{B}_{i}},K_{i})$
6:  Cluster $E_{\mathcal{B}_{i}}$ into $K^{*}$ clusters
7:  for $s\in\mathcal{B}_{i}$ do
8:   $c_{s}=\text{Consistency}(s)$
9:   $u_{s}=\text{Uniqueness}(s)$
10:   $w_{s}=\sigma\cdot c_{s}+(1-\sigma)\cdot u_{s}$
11:   for $f_{t}\in s$ do
12:     $w_{t}=w_{s}$
13:   end for
14:  end for
15:end for
16:return $\{w_{1},\dots,w_{T}\}$
@@
1:Input: Video frames $F_{\text{Video}}=\{f_{1},\dots,f_{T}\}$; Optional user query $Q_{\text{user}}$
2:Output: Frame-level importance scores
3:Initialize $\tilde{\mathcal{S}}\leftarrow\emptyset$
4:// Detect scene boundaries (Section 3.1)
@@
6:// Generate descriptions (Section 3.2)
[…]
8:// Score scenes using LLM (Section 3.3)
[…]
10:// Normalize scene scores and smooth frame-level scores (Section 3.4.1)
[…]
12:// Dynamically select segment duration and frame consistency weight
[…]
14:// Compute frame weights (Section 3.4.2)
[…]
16:for $i=1,\dots,N$ do
17:  Get scene boundaries $\mathcal{B}_{i}=[\text{start}_{i},\text{end}_{i}]$
18:  $\tilde{\mathcal{S}}_{i}\leftarrow\emptyset$
19:  for $t\in\mathcal{B}_{i}$ do
20:   // Compute final frame score
21:   $\tilde{f_{t}}\leftarrow F_{\text{smooth}}[t]\times F_{\text{weights}}[t]$
22:   Append $\tilde{f_{t}}$ to $\tilde{\mathcal{S}}_{i}$
23:  end for
24:  Normalize within scene: $\tilde{\mathcal{S}}_{i}\leftarrow F_{\text{norm}}(\tilde{\mathcal{S}}_{i})$
25:  Append $\tilde{\mathcal{S}}_{i}$ to $\tilde{\mathcal{S}}$
26:end for
27:return Frame-level importance scores $\tilde{\mathcal{S}}$
A.3 Scene detection full flow illustration and example
Figure 11 presents an additional pipeline flow example for the Bearpark_climbing video from the SumMe dataset . This example complements the general pipeline in Figure 2 in the main paper, and illustrates how the first stages—Selecting optimal threshold, Initial scene detection, and refined detected scene boundaries—is applied in practice. Figure 11(a) presents the flow procedure for partitioning the input set of frames into scenes based on a given threshold $\tau$.
(a) Scene boundary detection flow diagram.
(b) Determining the optimal threshold.
(c) Initial detected scenes based on frame intensity differences.
(d) Refined scene boundaries after merging short scenes using cosine similarity over mean scene embeddings.
Figure 11: Scene boundary detection and pipeline example for the Bearpark_climbing video (SumMe dataset). (a) Scene boundary flow diagram. (b) Threshold determination. (c) Initial detection. (d) Refined boundaries.
A.4 Dataset Details
The SumMe dataset comprises 25 videos (1–6 minutes each), covering diverse content such as holidays, events, and sports from both first-person and third-person perspectives. Each video was annotated by 15–18 users who provided key-fragment summaries, resulting in multiple user-generated summaries typically ranging from 5
The TVSum dataset includes 50 YouTube videos grouped into 10 categories (e.g., changing a tire, making a sandwich, visiting an award-winning shop), with 5 videos per category. Each video was annotated by 20 users who rated frame-level importance on a 1–5 scale [15].
The QFVS dataset [31] is built upon the UT Egocentric (UTE) dataset [30], consisting of four egocentric videos ranging from 3 to 5 hours. Each video is segmented into non-overlapping 5-second shots and annotated by three users with textual tags describing the visible content (e.g., CAR, ROAD, SKY, STREET). Each video is also associated with 45–46 user queries and corresponding user-annotated summaries. The user captions were aggregated per video-query into dense annotations, from which oracle (query-conditioned) summaries were derived using a greedy selection algorithm. These oracle summaries yield higher inter-user agreement compared to individual annotations.
A.5 Implementation details
We employ frame value density with a dynamic threshold (Section 3.1.1) for scene detection, using the contentDetector function from PySceneDetect [37] to identify scene boundaries. Frame embeddings are extracted using the CLIP/ViT-Large-Patch14 model [39], capturing both visual and contextual information effectively. For description generation, we use the LLAVA/Qwen-2.0-7B Video Language Model [54, 55] to produce textual representations at both the video and scene levels.
Memory management details. Experiments were conducted using approximately 12 GB of GPU memory. Sampled frames were divided into batches of 80 frames each. At a sampling rate of 1 FPS for full videos and 2 FPS for individual scenes, each batch corresponds to roughly 80 seconds and 40 seconds of video content, respectively. This configuration ensures detailed and temporally coherent descriptions while keeping memory usage within limits.
Optimizations for Longer Videos. For datasets like QFVS, where videos can exceed 5 hours with over 220k frames and multiple queries per video, processing each video–query pair individually is computationally prohibitive. To reduce overhead, we process all queries simultaneously for each scene by passing them together to the LLM, which assigns relevance scores for all queries in a single pass. This strategy can reduce processing time and resource usage by up to 50×.
(a) SumMe
(b) TVSum
Figure 12: Ablating different Image encoders and Visual context input for description generation. F1 scores results on the test set of SumMe dataset employing MinMax normalization and TVSum dataset employing Exponential (Exp) normalization across more diverse segment durations and $\sigma$ values ($\sigma\in[0.0,1.0]$).
Appendix B Additional experiments and results
B.1 Broader experiments
We conducted broader experiments to further validate our design choices. In Figures 12(a) and 12(b), we evaluated a more diverse set of segment durations compared to the experiments in Section  4.4(main paper). Specifically, we tested segment durations $W=[1,2,3,4]$, using different image encoders (DINO and CLIP) and varying the type of visual context input for description generation (Local or Global context). As shown, the previously identified best configurations (see Section 4.4 (main paper)) remain effective across both datasets. TVSum achieves the best performance with short segment durations and high frame weight values, while SumMe performs best with intermediate segment durations and moderate frame weight values. Notably, as previously mentioned, SumMe also performs competitively under the same configuration that works best for TVSum.
Table 4: Impact of video abstraction on summarization performance. Videos are categorized as Concrete or Abstract based on the temporal localization and visual saliency of key events.
Dataset Category #Videos Mean F1-score (%)
SumMe Concrete 11/25 49.0
Abstract 14/25 50.8
TVSum Concrete 36/50 57.5
Abstract 14/50 61.2
Video abstraction impact. We evaluated videos categorized as Concrete or Abstract based on the temporal localization of key events and their visual saliency, to assess whether our summarization pipeline is affected by the level of visual concreteness. As shown in Table 4, the mean F1-score is comparable—or slightly higher ($\Delta F_{1}<4\%$)—for Abstract videos, indicating that our pipeline is robust even when summarizing videos with more abstract or loosely defined events.
(a) SumMe
(b) TVSum
Figure 13: Ablating scene scores and within scene frame normalization. F1 scores results on the test set of SumMe dataset and TVSum dataset across more diverse segment durations ($W\in[1,2,3,4]$ seconds) and $\sigma$ values ($\sigma\in[0.0,1.0]$).
Normalization Impact. To evaluate the importance of applying normalization functions to the scene scores obtained from the LLM (see Section 3.3 (main paper)) and to the frame scores within each scene (see Section 3.4.3 (main paper)), we conducted experiments across a wider range of segment durations. Specifically, we applied Min-Max normalization for the SumMe dataset and Exponential normalization for the TVSum dataset, and compared performance with and without normalization.
As shown in Figures 13(a) and 13(b), across both datasets and for varying segment durations and frame weight values $\sigma$, applying a normalization function consistently outperforms not using one.
(a) SumMe
(b) TVSum
Figure 14: Ablating temporal smoothing function. F1 scores results on the test set of SumMe dataset and TVSum dataset across more diverse segment durations ($W\in[1,2,3,4]$ seconds) and $\sigma$ values ($\sigma\in[0.0,1.0]$).
Temporal smoothing function affect. Another key component in our pipeline is the temporal smoothing function. In Figure 14, we analyze the effect of employing this function versus not using it. It is clear that the smoothing function is critical, as its usage consistently yields better accuracy across all variations. Notably, its impact is more pronounced in the SumMe dataset (see Figure 14(a)). We suspect this is due to the misalignment between user-selected key-fragments and the segments produced by KTS, which the smoothing function helps to mitigate.
B.2 Prompt sensitivity
Large Language Models (LLMs) are known to be highly sensitive to prompt design. Small variations in phrasing or structure can lead to significantly different outputs, particularly in tasks involving subjective interpretation, such as video summarization. To investigate this sensitivity, we conducted experiments (see Figure 15) using four types of prompts:
  1. 1.
    Dummy prompt: A minimal placeholder prompt with no additional instructions or rubric for scoring scene importance.
  2. 2.
    Basic prompt: Provides the model with simple instructions to score scene importance based on overall relevance. Includes a rubric scale (1–100, with defined ranges for minimal to high importance) to guide scoring. We build upon the dummy prompt by adding the green parts (see Section A.1 in the appendix).
  3. 3.
    Instructive prompt: Builds upon the Basic prompt by adding more details about each score range and explicitly stating the goal of the task. This helps the model understand how each scene contributes to the overall summary and clarifies what each score represents. We build upon the basic prompt by adding the blue parts (see Section A.1 in the appendix).
  4. 4.
    Highly instructive prompt: An enhanced prompt containing explicit instructions on how to distribute importance scores across frames. Specifically, it guides the model to avoid a uniform distribution and instead assign higher scores to a small subset of key scenes. It maintains a mean score of approximately 50 out of 100, representing a balanced average importance level. We build upon the instructive prompt by adding the red parts (see Section A.1 in the appendix).
(a) SumMe
(b) TVSum
Figure 15: Ablating prompt sensitivity. F1 scores results on the test set of SumMe dataset and TVSum dataset across diverse segment durations ($W\in[1,3]$ seconds) and $\sigma$ values ($\sigma\in[0.0,1.0]$).
We evaluated the four prompt stages—Dummy, Basic, Instructive, and Highly Instructive—on the SumMe and TVSum datasets, using GPT-4o as the LLM judge. Experiments were conducted across different segment durations $W\in{1,3}$ and values of $\sigma$, which controls the consistency weighting of frames within each segment. The results, shown in Figure 15, demonstrate a clear performance trend as prompt specificity increases. In particular, the Highly Instructive prompt consistently outperforms the simpler prompt variants, indicating that explicitly constraining score distribution—by encouraging high importance scores for only a small subset of critical scenes while maintaining an overall mean near 50/100—enables the model to better approximate human-like importance judgments.
The improvement is more noticeable on the SumMe dataset (see Figure 15(a)), which is expected since SumMe represents a more challenging evaluation setting. Specifically, SumMe evaluates performance using the maximum F1 score across multiple annotators, effectively comparing the generated summary to the most similar individual human summary rather than an averaged consensus. This increases sensitivity to individual-level importance judgments and amplifies the impact of prompt design, making refined and structured prompts particularly beneficial. In contrast, TVSum averages annotations before evaluation, reducing variability across annotators and yielding a more stable, consensus-based reference. As a result, prompt-induced performance differences are less pronounced on TVSum dataset (see Figure 15(b)).
Table 5: Prompt sensitivity analysis on SumMe and TVSum datasets.